Grade 12th passPhysical Chemistry

a silver coin weight 2.4g is disolved in excess of HNO3 and then further treated with HCL sol in excess.The weight of AgCl formed is 2.87g.The % of Ag in coin is

Sueaj
8 Years agoGrade 12th pass

1 Answer

Swarnava Bag
ApprovedApproved Tutor Answer8 Years ago
No. Of moles of AgCl =No. Of moles of Ag in silver coin ,No. Of moles of AgCl=Wt.of AgCl formed÷Molecular weight of AgCl=2.87÷143.5=0.02 No. Of moles of Ag is 0.02 wt is 0.02*108=2.16 Wt %of Ag is=(2.14÷2.4)*100=90%